An aqueous solution of H 2 (CO 3 ) is added to an aqueous solution of Li(OH) and a reaction occurs. The

concentration of (OH) - is 3.75x10 -5 M. Calculate the concentration of the H 3 O +1 ions. Is the solution

more acidic or basic?


1
Expert's answer
2022-06-10T16:17:03-0400

pH = -log [H3O+]

pOH = -log [OH-]

in aqueous solution pH + pOH = 14


Using the above data, we find first the pOH, then the pH and then the H3O+ concentration.

pOH = -log (3,75*10-5) = 5- log(3,75) = 4,426


pH = 14 - pOH = 14 - 4,426 = 9,574


pH = -log [H3O+] = 9,574

[H3O+] = 2,667 * 10-10 M

If the pH is higher than 7, the solution is alkaline; if it is less than 7, it is acidic; if it is equal to 7, it is neutral.


Answer:

[H3O+] = * 10-10

solution is a more basic